# Self-organizing Multiagent Target Enclosing under Limited Information and Safety Guarantees

### v1

@@ Abstract
This paper introduces an approach to address the target enclosing problem using non-holonomic multiagent systems, where agents autonomously self-organize themselves in the desired formation around a fixed target. Our approach combines global enclosing behavior and local collision avoidance mechanisms by devising a novel potential function and sliding manifold. In our approach, agents independently move toward the desired enclosing geometry when apart and activate the collision avoidance mechanism when a collision is imminent, thereby guaranteeing inter-agent safety. We rigorously show that an agent does not need to ensure safety with every other agent and put forth a concept of the nearest colliding agent (for any arbitrary agent) with whom ensuring safety is sufficient to avoid collisions in the entire swarm. The proposed control eliminates the need for a fixed or pre-established agent arrangement around the target and requires only relative information between an agent and the target. This makes our design particularly appealing for scenarios with limited global information, hence significantly reducing communication requirements. We finally present simulation results to vindicate the efficacy of the proposed method.
@@
Multiagent Systems, Target Enclosing, Collision Avoidance, Motion Planning, Safety, Self-organizing Behavior.
@@ I Introduction
In recent times, autonomous vehicles have gained significant prominence in a range of tasks, including reconnaissance, surveillance, crop and forest monitoring, and search and rescue missions, e.g., [1, 2, 3, 4, 5]. Most of these tasks share a common behavior where an autonomous vehicle, known as the pursuer, monitors an object of interest, known as the target. In general, the pursuer’s stable motion around the target by maintaining the desired separation from the target is referred to as target enclosing [6] or encirclement/circumnavigation/entrapment when the enclosing orbit is a circle [7].
@@
One of the earliest approaches for target circumnavigation entails pursuers confining the target in a specific geometric shape posing as a formation control problem. The authors in [8, 9] utilize a cyclic pursuit strategy to cooperatively enclose a target, achieving and maintaining uniform angular spacing among the agents. Vector field guidance is another popular approach, where vector fields can be designed to create a limit cycle or periodic orbit around the target, guiding agents to desired geometric shapes [10, 11]. However, each orbit may require a separate field, thus limiting its applicability in a multiagent scenario. Several other guidance strategies have also been developed for trapping targets in circular orbits, based on differential geometry [12], model predictive control [13], and reinforcement learning [14]. Many of the works in the target enclosing have focused on reducing the information required for developing the control laws. For example, some strategies utilize only range-only measurements [15, 16, 17], bearing angle measurement [18, 19], range and bearing information [20, 7, 21]. The general moving target enclosing problem using a single agent was studied in [22, 23], whereas the authors in [24] guided multiple agents to the desired proximity with equal angular spacing. Note that these works did not account for explicit inter-agent safety.
Most studies on target enclosing have focused on ensuring that pursuers reach the target without considering their safety or coordinating their movements with other pursuers. However, this approach can result in collisions between pursuers, which may result in mission failure. To address this issue, some studies have introduced decentralized methods for collision avoidance between pursuers, allowing them to orbit the target while maintaining a safe distance from each other [8, 9, 13, 24, 25, 26]. However, these methods often rely on strict formations, such as equal angular spacing, which can limit the pursuers’ movement and increase the complexity of the guidance strategy. By eliminating the need for rigid formations and reducing the number of agents that each pursuer must communicate with, we can simplify the guidance strategy and improve its robustness to changes in the number of pursuers in the enclosing formation.
Motivated by the aforementioned studies, our work aims to create efficient target-enclosing guidance laws that do not require extensive information exchange or predefined agent configurations. Our main contributions include developing a distributed guidance law for pursuers to safely enclose a target while avoiding collisions, using a minimal amount of information, and no fixed agent formations. Each pursuer only coordinates with one other pursuer when necessary to prevent collisions, thereby ensuring inter-agent safety. Our approach introduces a novel decision-making protocol for selecting the appropriate pursuer for coordination and designing a guidance strategy to ensure collision avoidance. We design the guidance law based on a custom potential function that combines attractive and repulsive forces to guide the pursuers toward the desired orbit around the target safely and efficiently. We also demonstrate that our proposed guidance law is robust to the motion of other pursuers and requires only relative information, potentially eliminating the need for costly agent-to-agent communication.
@@ II-A Kinematics of Agents’ Relative Motions
Consider a multiagent system consisting of $n$ pursuers and a single stationary target $T$. We use $\mathcal{L}$ to denote the set of the pursuer, that is, =$\mathcal{L}=\{P_{1},P_{2},\ldots,P_{n}\}$. Figure 1 depicts the engagement geometry between the pursuers and the target, where $[x_{i},y_{i}]^{\top}\in\mathbb{R}^{2}$ denotes the position of the $i$th pursuer in the inertial frame of reference, and $\chi_{i}\in[-\pi,\pi)$ denotes its heading angle. Assuming the vehicles to be point masses, the motion of the $i$th pursuer is governed by
@@
Figure 1: Target enclosing scenario in the inertial frame of reference.
The engagement geometry between the vehicles in the relative frame of reference is depicted in Figure 2, where $r_{ij}$ denotes the distance between the $i$th and the $j$th pursuer ($i\neq j$) and accordingly $r_{iT}$ signifies the distance with the target. Similarly, $\theta_{iT}$ denotes the line-of-sight (LOS) angle for the $i$th pursuer with respect to the target and $\theta_{ij}$ is the LOS angle between the $i$th and the $j$th pursuer referenced from the $i$th pursuer. The bearing angle of the $i$th pursuer with respect to the target is $\sigma_{i}=\chi_{i}-\theta_{iT}$. Thus, the kinematics of the relative motion for the multiagent system is given by
@@
These equations are obtained by projecting the $i$th pursuer’s relative velocities in the directions parallel/normal to the $i$th pursuer-target LOS.
Figure 2: Target enclosing scenario in the relative frame of reference.
@@ II-B Direction/Sense of Target Enclosing
For the target-centric relative system in (II-A), $r_{iT}$ is measured from $P_{i}$ to target and $\theta_{iT}$ is measured in a counter-clockwise sense from the horizontal reference. Therefore, positive angular rates pertain to the counter-clockwise revolution of the pursuers about the target. Further, it is evident from (1b) that $\dot{\theta}_{iT}<0,\;\forall\;\sigma_{i}\in(0,\pi)$, $\dot{\theta}_{iT}>0,\forall\,\sigma_{i}\in(-\pi,0)$ and $\dot{\theta}_{iT}=0,\forall\,\sigma_{i}\in\{0,\pm\pi\}$.
@@
$d_{ij}$ $=r_{iT}-r_{jT},$ (2a)
$\psi_{ij}$ $=\theta_{iT}-\theta_{jT},$ (2b)
where $\psi_{ij}\in[-\pi,\pi)$, as depicted in Figure 3, where the agents (the pursuers) are enclosing the target in a clockwise direction. In Figure 3, the dotted circles represent the loiter circles around the target with a radius equal to the instantaneous pursuer-target distance. The positive values of $d_{ij}$ imply that the $i$th pursuer is on a larger loiter circle compared to the $j$th pursuer. Hence, $d_{ij}>0$ implies $\dfrac{{r}_{iT}}{r_{jT}}>1$, and $d_{ij}<0$ implies that $10<\dfrac{{r}_{iT}}{r_{jT}}<$ 1. Similarly, the positive values of $\psi_{ij}$ result in a scenario when $\theta_{jT}$ leads $\theta_{iT}$ in a clockwise sense. The red arrows in Figure 3 denote the positive conventions of $d_{ij}$ and $\psi_{ij}$ considered in this paper. It is also worth noting that for any two agents $P_{i}$ and $P_{j}$, $d_{ij}$ and $\psi_{ij}$ together represent the inter-agent radial distance $r_{ij}$. From Figure 2, using the cosine angle formula from triangle $P_{i}-P_{j}-T$, we can obtain the distance between $P_{i}$ and $P_{j}$ as, $r_{ij}^{2}=r_{iT}^{2}+r_{jT}^{2}-2r_{iT}r_{jT}\cos\psi_{ij}$, which reduces to,
@@
on using $r_{iT}=r_{jT}+d_{ij}$ from (2a). Hence, if both $d_{ij}\to 0$ (coincident loiter circles) and $\psi_{ij}\to 0$ (coincident pursuer-target LOS), then it essentially implies that $r_{ij}\to 0$, which means inter-agent collision since more than one pursuer occupies the same position.
@@ II-C Design of Control Objectives
Figure 4: Engagement between pursuers without safety considerations.
To elucidate the implications of Remark 2, let us analyze a typical case of target enclosing involving two pursuers, as shown in Figure 4. If $P_{j}$ is initially located closer to the target than $P_{i}$ such that $d_{ij}>0$ and $\psi_{ij}<0$, then $\dot{\theta}_{iT}<\dot{\theta}_{jT}$. This essentially means that the component of the relative velocity of $P_{j}$ normal to $r_{jT}$ is greater than that of $P_{i}$, thereby making $\dot{\psi}_{ij}>0$. This leads to coincident pursuer-target LOS between $P_{i}$ and $P_{j}$ at some instant of time during target enclosing. Similarly, the component of the relative velocity of $P_{j}$ along $r_{jT}$ will be less than that of $P_{i}$, thereby making $\dot{d}_{ij}<0$. This results in the pursuers being on the same loiter circle at some instant of time during target enclosing. Consequently, a collision between $P_{i}$ and $P_{j}$ may eventually occur, as demonstrated in Figure 4, where ${(\cdot)}^{{}^{\prime}}$ and ${(\cdot)}^{{}^{\prime\prime}}$ represent the variables related to the pursuers as time progresses since both $d_{ij}\to 0$ and $\psi_{ij}\to 0$ as a result of $\dot{d}_{ij}<0$ and $\dot{\psi}_{ij}>0$. Therefore, one may speculate that inter-agent collisions can be prevented if such conditions/scenarios never arise during target enclosing.
The previous analysis showed that collisions between pursuers occur only under specific circumstances. Our method takes advantage of this fact to restrict each pursuer’s motion to avoid collisions solely with those pursuers that have the potential to collide in the future. It is important to note that when there are multiple pursuers, several of them may collide with $P_{i}$. Consequently, we devise strategies to prevent collisions with a single pursuer while taking into account the potential collisions with all the other pursuers. To this end, we first identify the set of all pursuers that are within the sensing range of the $i$th pursuer (i.e., $r_{ij}\leq r_{s}$) and may potentially collide with it. Let us refer to this set as
$\mathcal{N}_{i}\coloneqq\left\{P_{j}\in\mathcal{L}\,\big{|}\ d_{ij}>0\,\&\,\psi_{ij}<0\right\},~{}\forall\,i\neq j,$ (5)

### v2

@@ Abstract
This paper introduces an approach to address the target enclosing problem using non-holonomic multiagent systems, where agents self-organize on the enclosing shape around a fixed target. In our approach, agents independently move toward the desired enclosing geometry when apart and activate the collision avoidance mechanism when a collision is imminent, thereby guaranteeing inter-agent safety. Our approach combines global enclosing behavior and local collision avoidance mechanisms by devising a special potential function and sliding manifold. We rigorously show that an agent does not need to ensure safety with every other agent and put forth a concept of the nearest colliding agent (for any arbitrary agent) with whom ensuring safety is sufficient to avoid collisions in the entire swarm. The proposed control eliminates the need for a fixed or pre-established agent arrangement around the target and requires only relative information between an agent and the target. This makes our design particularly appealing for scenarios with limited global information, hence significantly reducing communication requirements. We finally present simulation results to vindicate the efficacy of the proposed method.
@@
Multiagent Systems, Target Enclosing, Collision Avoidance, Motion Planning, Safety, Self-organization.
@@ I Introduction
Autonomous vehicles have gained significant prominence in a range of tasks, including reconnaissance, surveillance, crop and forest monitoring, and search and rescue missions, e.g., [1, 2, 3, 4, 5]. Most of these tasks share a common behavior where an autonomous vehicle, known as the pursuer, monitors an object of interest, known as the target. In general, the pursuer’s stable motion around the target by maintaining the desired separation from the target is referred to as target enclosing [6] or encirclement/circumnavigation/entrapment when the enclosing orbit is a circle [7].
@@
One of the earliest approaches for target circumnavigation entails pursuers confining the target in a specific geometric shape posing as a formation control problem. The authors in [8, 9] utilize a cyclic pursuit strategy to cooperatively enclose a target, achieving and maintaining uniform angular spacing among the agents. Vector field guidance is another popular approach, where vector fields can be designed to create a limit cycle or periodic orbit around the target, guiding agents to desired geometric shapes [10, 11]. However, each orbit may require a separate field, thus limiting its applicability in a multiagent scenario. Several other guidance strategies have also been developed for trapping targets in circular orbits, based on differential geometry [12], model predictive control [13], and reinforcement learning [14]. Many of the works in the target enclosing have focused on reducing the information required for developing the control laws. For example, some strategies utilize only range-only measurements [15, 16, 17], bearing angle measurement [18, 19, 20], range and bearing information [21, 7, 22]. In [23, 24], authors developed a flexible guidance strategy that significantly enhances the pursuer’s maneuverability by allowing it to enclose the target in arbitrary shapes.
While most of the aforementioned studies focus on developing guidance strategies for a single pursuer, many researchers have explored cooperative targets enclosing [8, 9, 25, 26, 27]. These approaches aim to reduce the complexity of guidance strategies by requiring every pursuer to coordinate with only some of the neighboring agents. The authors in [8, 9] designed a guidance strategy that arranges the pursuers in a circular formation, achieving uniform angular spacing on the orbit around the target. The works in [25, 26, 27], utilized various rigid formations between the pursuers and the target converging to desired relative configurations on the enclosing shape. Note that most of the above-mentioned works did not account for explicit inter-agent safety and focused on guiding the pursuers to the enclosing geometry without considering the possibility of collisions. Although a rigid formation shape may inherently ensure that no two pursuers occupy the same position at the steady state, additional considerations are necessary to prevent collisions at all times.
Therefore, recent works have incorporated collision-avoidance guarantees in target-enclosing strategies [28, 29, 30, 31, 32, 33], steering the pursuers on a collision-free trajectory towards the enclosing shape around the target. For example, in [28, 29, 31], pursuers achieve uniform angular spacing, while [30, 32] allow for predefined uneven angular spacing among pursuers on a fixed orbit around the target. The guidance strategy in [33] not only avoids collisions among pursuers but also with static obstacles. Note that all the above works on multi-agent target enclosing utilize a rigid formation structure that significantly limits each pursuer’s maneuverability and necessitates vehicle-to-vehicle communication. Each pursuer must simultaneously meet formation objectives (desired position and orientation) relative to both the target and multiple other pursuers to maintain the desired formation shape. In practice, pursuers may not be able to accurately track the desired position due to external disturbances or steady-state errors. Therefore, a strict formation structure may not be suitable for multi-agent target enclosing, as it constrains the pursuers to maneuver in accordance with multiple pursuers, and the failure of even a single pursuer may be catastrophic for the entire formation.
This work is motivated by the need to develop safe motion coordination strategies for target enclosing that do not rely on rigid formation structures. Traditional target enclosing strategies can be overly restrictive, limiting the maneuverability of individual pursuers and requiring constant communication to maintain precise positions. This rigidity not only increases the computational and communication burden on the system but also reduces its robustness and adaptability in dynamic environments. The proposed flexible coordination strategy offers significant advantages as it allows pursuers to maneuver more freely, adapting to changing conditions and unexpected obstacles (in the form of neighboring pursuers) without compromising the overall goal of the target enclosing. By reducing the adherence to strict formation maintenance, the proposed strategy also minimizes communication efforts and the need for extensive prior information about the target and the environment. This can lead to more efficient and scalable solutions, particularly in scenarios where the environment is contested or the communication infrastructure is limited. The contributions of this paper can be summarized as follows:
  1. 1.
    We develop a distributed guidance law for the pursuers to safely enclose a stationary target, allowing them to autonomously arrange themselves on the enclosing orbit. This work builds on our previous research [6, 7, 15, 22, 23, 24], where we developed flexible enclosing strategies for a single pursuer-target pair. We extend those strategies to accommodate multiple pursuers, ensuring explicit safety guarantees. Unlike previous methods on multi-agent target enclosing [8, 9, 25, 26, 27, 28, 29, 30, 31, 32, 33] that rely on fixed or predetermined formations that restricts the final configuration of pursuers about the target, our approach enables self-organizing behavior. Such a behavior inherently allows the enclosing formation to adapt to a changing number of pursuers, with vehicles ensuring a safe distance among them.
  2. 2.
    Our approach steers the pursuers to arbitrary points on the enclosing orbit and ensures they do not collide with each other. This allows them to achieve cooperative target enclosing with any final relative configuration and eliminates the need for multiple pursuers to coordinate maneuvers to avoid collisions. Consequently, our method significantly enhances the maneuverability of each pursuer compared to previous works [8, 9, 25, 26, 27, 28, 29, 30, 31, 32, 33], which rely on fixed formation structures that constrain each pursuer’s path to maintain a strict formation. This approach differs from most of our previous work [6, 7, 15, 22, 23, 24], which, although flexible, did not account for vehicle safety explicitly in changing operating conditions.
  3. 3.
    We introduce a novel decision-making protocol for selecting the appropriate pursuer for coordination and designing a collision avoidance strategy for inter-agent safety. This protocol ensures that each pursuer only coordinates with at most one other pursuer when necessary to prevent collisions. Unlike previous approaches [28, 29, 30, 31, 32, 33] that require extensive coordination and communication between multiple pursuers, our strategy significantly reduces the complexity of control design, preventing the unnecessary maneuvering of multiple vehicles in order to avoid collisions, which may also help in reducing the overall control effort.
  4. 4.
    The proposed guidance laws are developed utilizing custom potential function that combines attractive forces with switching repulsive forces to guide the pursuers toward the desired orbit around the target safely and efficiently. This is unlike previous potential function-based target enclosing approaches [10, 11, 34, 35] that utilize swarming or flocking-based methods (where repulsive force acts continuously among all neighboring pursuers). Our approach activates repulsive forces only when necessary and only towards one other neighbouring pursuer enabling more independent maneuvers with reduced information.
  5. 5.
    We demonstrate that our proposed guidance law is robust to the motion of other pursuers in the swarm and requires only relative information. Unlike previous approaches (including some of our previous works [1, 15]) that rely on global information or information on other pursuers’ maneuvers to achieve collision-free target enclosing, our approach potentially eliminates the need for costly agent-to-agent communication and can perform efficiently in environments where global information is not readily available.
@@ II-A Kinematics of Agents’ Relative Motions
Consider a multiagent system consisting of $n$ pursuers and a single stationary target $T$. We use $\mathcal{L}$ to denote the set of the pursuer, that is, =$\mathcal{L}=\{P_{1},P_{2},\ldots,P_{n}\}$. Figure 1(a) depicts the engagement geometry between the pursuers and the target, where $[x_{i},y_{i}]^{\top}\in\mathbb{R}^{2}$ denotes the position of the $i$th pursuer in the inertial frame of reference, and $\chi_{i}\in[-\pi,\pi)$ denotes its heading angle. Assuming the vehicles to be point masses, the motion of the $i$th pursuer is governed by
@@
The engagement geometry between the vehicles in the relative frame of reference is depicted in Figure 1(b), where $r_{ij}$ denotes the distance between the $i$th and the $j$th pursuer ($i\neq j$) and accordingly $r_{iT}$ signifies the distance with the target. Similarly, $\theta_{iT}$ denotes the line-of-sight (LOS) angle for the $i$th pursuer with respect to the target and $\theta_{ij}$ is the LOS angle between the $i$th and the $j$th pursuer referenced from the $i$th pursuer. The bearing angle of the $i$th pursuer with respect to the target is $\sigma_{i}=\chi_{i}-\theta_{iT}$. Thus, the kinematics of the relative motion for the multiagent system is given by
@@
These equations are obtained by projecting the $i$th pursuer’s relative velocities in the directions parallel/normal to the $i$th pursuer-target LOS. Assuming the pursuers have onboard sensors, we utilize the relative kinematics to develop the control laws minimizing the need for vehicle-to-vehicle communication. Consequently, other pursuers are positioned within each pursuer’s sensing range $r_{S}$ to be detected.
(a) Inertial frame of reference.
(b) Relative frame of reference.
Figure 1: Multi-agent Target enclosing geometry.
@@ II-B Direction/Sense of Target Enclosing
For the target-centric relative system in (1), $r_{iT}$ is measured from $P_{i}$ to target and $\theta_{iT}$ is measured in a counter-clockwise sense from the horizontal reference. Therefore, positive angular rates pertain to the counter-clockwise revolution of the pursuers about the target. Further, it is evident from (1b) that $\dot{\theta}_{iT}<0,\;\forall\;\sigma_{i}\in(0,\pi)$, $\dot{\theta}_{iT}>0,\forall\,\sigma_{i}\in(-\pi,0)$ and $\dot{\theta}_{iT}=0,\forall\,\sigma_{i}\in\{0,\pm\pi\}$.
@@
$d_{ij}=r_{iT}-r_{jT},\;\;\psi_{ij}=\theta_{iT}-\theta_{jT},$ (2)
where $\psi_{ij}\in[-\pi,\pi)$, as depicted in Figure 2, where the agents (the pursuers) are enclosing the target in a clockwise direction. In Figure 2, the dotted circles represent the loiter circles around the target with a radius equal to the instantaneous pursuer-target distance. The positive values of $d_{ij}$ imply that the $i$th pursuer is on a larger loiter circle compared to the $j$th pursuer. Hence, $d_{ij}>0$ implies $\dfrac{{r}_{iT}}{r_{jT}}>1$, and $d_{ij}<0$ implies that $10<\dfrac{{r}_{iT}}{r_{jT}}<$ 1. Similarly, the positive values of $\psi_{ij}$ result in a scenario when $\theta_{jT}$ leads $\theta_{iT}$ in a clockwise sense. The red arrows in Figure 2 denote the positive conventions of $d_{ij}$ and $\psi_{ij}$ considered in this paper. It is also worth noting that for any two agents $P_{i}$ and $P_{j}$, $d_{ij}$ and $\psi_{ij}$ together represent the inter-agent radial distance $r_{ij}$. From Figure 1(b), using the cosine angle formula from triangle $P_{i}-P_{j}-T$, we can obtain the distance between $P_{i}$ and $P_{j}$ as, $r_{ij}^{2}=r_{iT}^{2}+r_{jT}^{2}-2r_{iT}r_{jT}\cos\psi_{ij}$, which reduces to,
@@
on using $r_{iT}=r_{jT}+d_{ij}$ from (2). Hence, if both $d_{ij}\to 0$ (coincident loiter circles) and $\psi_{ij}\to 0$ (coincident pursuer-target LOS), then it essentially implies that $r_{ij}\to 0$, which means inter-agent collision since more than one pursuer occupies the same position.
@@ II-C Design of Control Objectives
(a) Without safety.
(b) With safety .
Figure 3: Engagement geometry among pursuers in target-enclosing.
To elucidate the implications of Remark 2, let us analyze a typical case of target enclosing involving two pursuers, as shown in Figure 3(a). If $P_{j}$ is initially located closer to the target than $P_{i}$ such that $d_{ij}>0$ and $\psi_{ij}<0$, then $\dot{\theta}_{iT}<\dot{\theta}_{jT}$. This essentially means that the component of the relative velocity of $P_{j}$ normal to $r_{jT}$ is greater than that of $P_{i}$, thereby making $\dot{\psi}_{ij}>0$. This leads to coincident pursuer-target LOS between $P_{i}$ and $P_{j}$ at some instant of time during target enclosing. Similarly, the component of the relative velocity of $P_{j}$ along $r_{jT}$ will be less than that of $P_{i}$, thereby making $\dot{d}_{ij}<0$. This results in the pursuers being on the same loiter circle at some instant of time during target enclosing. Consequently, a collision between $P_{i}$ and $P_{j}$ may eventually occur, as demonstrated in Figure 3(a), where ${(\cdot)}^{{}^{\prime}}$ and ${(\cdot)}^{{}^{\prime\prime}}$ represent the variables related to the pursuers as time progresses since both $d_{ij}\to 0$ and $\psi_{ij}\to 0$ as a result of $\dot{d}_{ij}<0$ and $\dot{\psi}_{ij}>0$. Therefore, one may speculate that inter-agent collisions can be prevented if such conditions/scenarios never arise during target enclosing.
The previous analysis showed that collisions between pursuers occur only under specific circumstances. Our method takes advantage of this fact to restrict each pursuer’s motion to avoid collisions solely with those pursuers that have the potential to collide in the future. It is important to note that when there are multiple pursuers, several of them may collide with $P_{i}$. Consequently, we devise strategies to prevent collisions with a single pursuer while taking into account the potential collisions with all the other pursuers. To this end, we first identify the set of all pursuers that are within the sensing range $r_{s}$ of the $i$th pursuer (i.e., $r_{ij}\leq r_{s}$) and may potentially collide with it. Let us refer to this set as
$\mathcal{N}_{i}\coloneqq\left\{P_{j}\in\mathcal{L}\,\big{|}\ d_{ij}>0,\psi_{ij}<0\,\&\,r_{ij}\leq r_{s}\right\},~{}\forall\,i\neq j,$ (5)